2 and 1 half over 3 and 1 third = Y over 4 and 1 fourth. What is Y?
Answer by checkley77(12844)

2.5/3+1/3=y/4+1/4
(2.5+1)/3=(y+1)/4
3.5/3=(y+1)/4 cross multiply.
3(y+1)=4*3.5
3y+3=14
3y=14-3
3y=11
y=11/3 ans.
Proof:
2.5/3+1/3=(11/3)/4+1/4
3.5/3=11/3*1/4+1/4
3.5/3=11/12+1/4
3.5/3=(11+3)/12
3.5/3=14/12
3*14=3.5*12
42=42